OpenStudy (anonymous):

Assume that y varies inversely as x. If x = 12 when y = 9, what is x when y = -3? 108 -36 27 -4 Assume that y varies inversely with x. If y = 3.75 when x = 9, what is x when y = 6.75? 33.75 1.8 5 60.75 Use the Law of Levers to solve the problem, x1y122. The number of chairs in a row varies inversely as the distance between them. When they are 10 centimeters apart, 27 chairs fit in a row. If 30 chairs are needed in a row, how many centimeters apart should they be placed? 9 cm 12 cm 15 cm 8 cm

OpenStudy (vishal_kothari):

first one is -36...

OpenStudy (anonymous):

can you help me with the other two?

OpenStudy (anonymous):

yeah, -36

OpenStudy (vishal_kothari):

second one is 5

OpenStudy (vishal_kothari):

xy=k

OpenStudy (vishal_kothari):

as it is inverse...

OpenStudy (vishal_kothari):

k=3.75*9

OpenStudy (vishal_kothari):

thus, 3.75*9=6.75*x solving the eqn.. you get x=5

OpenStudy (anonymous):

Thanks! You dont knowthe other one? And hey are you russian

OpenStudy (vishal_kothari):

no i am indian....

OpenStudy (vishal_kothari):

are you russian?

OpenStudy (anonymous):

no haha i was just wondering because of your name.

OpenStudy (vishal_kothari):

can you write clearly the third one...

OpenStudy (vishal_kothari):

?